Questions Evaluate S 45ydo over the surface S, the first octant of the plane +2y+2z = 4. a) 16 b) 16 3 c) 16 3 d) O-16 8
Answer #1

Js Azy do so x+2y+22=4 iu first quadrant Il 4xydo Jls 4 xy J1H (22) + (zje dy de 2= 4 - 2 - 2 1 z = 2- x 2 - 4 az da 2 = -1 -|| 4 xy do 3/4 x ( 4-1² dx 2-0 4 4 SI Axy do = 3) x ( 16 +2² -8x) dx 4220 = 30 + (16x + 2²-82²) dx 4 xzo 4 + 4 [ 16 2 3 JS 4hence option a is correct.....